### v1

# Discovery of a nova super-remnant surrounding the recurrent nova LMCN 1971-08a in the Large Magellanic Cloud

Sternwartstraße 7, 96049 Bamberg, Germany 22institutetext: Astrophysics Research Institute, Liverpool John Moores University, Liverpool, L3 5RF, UK 33institutetext: NSF’s NOIRLab/CTIO, Casilla 603, La Serena, 1700000, Chile 44institutetext: School of Science, Western Sydney University, Locked Bag 1797, Penrith, NSW 2751, Australia 55institutetext: Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 66institutetext: Eureka Scientific, Inc. 2452 Delmer Street, Suite 100, Oakland, CA 94602-3017, USA
Discovery of a nova super-remnant surrounding the recurrent nova LMCN 1971-08a in the Large Magellanic Cloud
A nova super-remnant (NSR) is a greatly-extended structure grown by repeated nova eruptions sweeping the surrounding material away from a nova into a dense outer shell and are predicted to form around all novae. To date, four NSRs are known, with three located in the Galaxy and one residing in M31. Here we present the discovery of the first NSR in the Large Magellanic Cloud and only the second extragalactic nova shell to be identified, hosted by the recurrent nova LMCN 1971-08a. The structure is coincident with the nova, has a circular morphology and is visible in narrowband H and filters but very faint in , as expected. Hi data also potentially reveal the existence of a coincident structure. Further, with a diameter of pc, this NSR is the largest example yet found, with models indicating an shell expanding at into the surrounding medium and an age of 2.4 Myr. The existence of the NSR also suggests that LMCN 1971-08a may have a much shorter recurrence period than currently presumed.
Key Words.:
Stars: novae, cataclysmic variables – Stars: individual: LMCN 1971-08a – Magellanic Clouds – ISM: structure – Radio continuum: ISM1 Introduction
Recurrent novae (RNe) are observationally-defined as accreting white dwarf (WD) binaries that have exhibited more than one nova eruption. Such eruptions are driven by a thermonuclear runaway on the WD’s surface (Starrfield et al., 1972) that leads to a portion of the accreted material being ejected at velocities up to thousands of kilometers per second (O’Brien et al., 2001) that, in some cases, form expanding sub-parsec nova shells (Shara et al., 2024b; Santamaría et al., 2025). After the eruption, the resumption of accretion onto the WD (Worters et al., 2007) prepares the nova system for subsequent eruptions. Furthermore, accumulation of material retained on the WD following each eruption can grow the WD (Yaron et al., 2005; Hillman et al., 2016) to the Chandrasekhar limit to potentially explode as a type Ia supernova (Whelan & Iben, 1973).
Nova super-remnants (NSRs) are shell-like structures orders of magnitude larger than singular eruption nova shells (for a review see Healy-Kalesh, 2024). The first NSR (Darnley et al., 2019) was discovered in the Andromeda Galaxy around the rapidly recurring nova, M31N 2008-12a, and has been formed by the ejecta of repeated nova eruptions over millenia excavating tens of thousands of solar masses of surrounding interstellar medium from around the central binary into a high density shell (Darnley et al., 2019). Theoretically, such structures should encompass all novae (Healy-Kalesh et al., 2023) though their size on the sky and low surface brightnesses make NSRs difficult to find in optical data. However, with the capabilities of the recently deployed Condor Array Telescope (Lanzetta et al., 2023), optical NSRs have now been discovered around the Galactic RNe KT Eridani (Shara et al., 2024a; Healy-Kalesh et al., 2024c), T Coronae Borealis (Shara et al., 2024c) and RS Ophiuchi (Shara et al., 2025)111A possible cavity component of the NSR associated with RS Ophiuchi was identified in far-infrared data (Healy-Kalesh et al., 2024a).. These NSRs have diameters ranging from pc; have emission most prominent in H, and ; lack strong emission as a result of ionisation from low velocity shocks; and exhibit low expansion velocities in their outer shells. While the NSR around M31N 2008-12a appears to have a more well-defined elliptical shell as it is viewed pole on (Darnley et al., 2019), the Galactic NSRs have more patchy morphologies likely as a result of the NSR growing within the irregular interstellar medium (ISM) and possibly from being observed at different viewing angles. Furthermore, the NSRs surrounding T CrB (Shara et al., 2024c) and RS Oph (Shara et al., 2025) appear bilobed, possibly resulting from the ejecta being initially influenced by the accretion disk around the WD. A search for more NSRs in M31 and the Large Magellanic Cloud (LMC) did not lead to any more examples being found (Healy-Kalesh et al., 2024b) but this apparent dearth was attributed to the structures around other ‘younger’ systems being considerably fainter.
LMCN 1971-08a is one of the four currently known recurrent novae in the LMC. It exhibited eruptions in 1971 (Graham, 1971) and 2009 (Liller, 2009) and therefore has a recurrence period (or inter-eruption interval) of /N years (Bode et al., 2016), where N may be one or a small integer. While only a spectrum of the first eruption was obtained in 1971 with no further follow-up (Graham, 1971; Bode et al., 2016), the more recent eruption in 2009 was followed in great detail, being observed in the optical, UV, near-infrared (Bode et al., 2016) and X-rays (Bode et al., 2016; Orio et al., 2021). These observations revealed a very-fast declining nova comprising a WD accreting material at an elevated rate of from a sub-giant companion with a probable orbital period of 1.2 days (Bode et al., 2016). The ejecta velocities seen from the nova ranged from (Orio et al., 2009; Bode et al., 2016) with shocks likely present in the ejecta before day nine of the eruption (Bode et al., 2016).
In this Letter, we present the discovery of a nova super-remnant around the recurrent nova LMCN 1971-08a; the first nova super-remnant to be found in the LMC. In Section 2 we outline the datasets used in this work before describing the characteristics of the NSR across optical and radio data in Section 3. In Section 4 we place the discovery within context of the growing field of nova super-remnants and rule out other possible origins before summarising our findings in Section 5.
2 Optical and radio data
Within this study, several datasets were utilised across the optical and radio. The Magellanic Cloud Emission Line Survey (MCELS; Smith & Team, 1999) observed the central of the LMC with five filters including narrow-band H, and filters. The DeMCELS survey (Points et al., 2024) is a contemporary higher-resolution emission-line survey of the Magellanic Clouds undertaken with the Dark Energy Camera (DECam; Flaugher et al., 2015) narrow-band H and filters. The continuum-subtracted narrowband data from both MCELS and DeMCELS were utilised for this work. The HI4PI survey (HI4PI Collaboration et al., 2016) is an all-sky neutral atomic hydrogen (Hi) survey constructed from the Effelsberg-Bonn Hi Survey (EBHIS; Kerp et al., 2011) and the Galactic All-Sky Survey (GASS; McClure-Griffiths et al., 2009) data sets, thereby providing better angular resolution and sensitivity than previous all-sky Hi surveys. The MeerKAT radio data (Cotton et al. in prep.) used in this work are from the MeerKAT array (Jonas & MeerKAT Team, 2016), a new generation radio telescope with high resolution and sensitivity such that it is capable of resolving structures in the LMC (see, e.g. Sasaki et al., 2025).
3 LMCN 1971-08a nova super-remnant
A coherent, circular shell-like structure spatially coincident with LMCN 1971-08a appears in the MCELS H data (see Figure 1), with the same shell visible in the MCELS image but negligible in MCELS , as is expected for NSRs (Darnley et al., 2019; Shara et al., 2024a, c, 2025). While clearly evident in the MCELS data, the high resolution DeMCELS H and imaging of the NSR emphasizes the structure further, as shown in Figure 2 (with the approximate extent of the NSR indicated).
With a distance of kpc to the LMC (Pietrzyński et al., 2019), the NSR has a projected diameter of approximately 200 pc. As seen in the optical data of Figure 1, the NSR is brighter to the northeast and southwest with a fainter boundary connecting the two components to the northwest, defining the NSR outer shell edge. The shell thickness is taken to be from the southwest edge as the inner and outer boundaries of this part of the NSR shell are pc and pc from the nova, respectively. This possible compression front may indicate, as with a similar feature in the KT Eri NSR, interaction of a bow shock with an earlier phase of the remnant’s evolution (Healy-Kalesh et al., 2024c). Additionally, there appears to be an apparent interaction with a complex structure to the west (Figure 3) composed of many Hii regions (Pellegrini et al., 2012).
Following Points et al. (2024), the H and surface brightnesses of the bright northeast component of the NSR are and , respectively. The H and surface brightnesses of the bright southwest component are marginally fainter radiating at and , respectively.
3.1 NSR in radio data
As shown in Figure 1, within the Hi data cube from the HI4PI survey (HI4PI Collaboration et al., 2016) sliced at and , there does appear to be correlated structure coincident with the H shell. Specifically, there appears to be a higher level of emission from the southwestern portion of the NSR in the slice and a possible shell boundary to the northeast in the slice. A velocity distribution across the NSR with a range of in the declination and right ascension direction (see Figure 4 for the slits used) does reveal an arc with a range of as shown in Figure 5, possibly indicating an outer shell expansion velocity of . Within MeerKAT radio data (Cotton et al. in prep.), there does not strongly appear to be a corresponding structure, as shown in Figure 1. By estimating the RMS noise level in the region of the NSR shell, we derive an upper limit of Jy / beam.
3.2 Simple model of NSR evolution
To understand the evolutionary history of the NSR, we employed the hydrodynamical code Morpheus (Vaytet et al., 2007) following Healy-Kalesh et al. (2023, 2024c), to produce a simple model of the growth of such a remnant with a radius of pc. For the eruption characteristics of LMCN 1971-08a, we utilised the data from (Bode et al., 2016), taking the recurrence period of the nova to be 38 years; the mass loss phase to be 10.4 days (equal to the band decline time); the ejecta velocity to be ; and the ejecta mass to be .
To estimate the surrounding ISM density for the model, we took the Hi data velocity cube from the HI4PI survey (HI4PI Collaboration et al., 2016), converted the integrated intensity in the range of into column density (using equation 2 from HI4PI Collaboration et al., 2016) and found the average within of the nova (to incorporate the NSR and its surroundings). Assuming a depth of 5 kpc (estimated from the NSR’s location in Figure 9 of Subramanian & Subramaniam, 2009) gives an ISM density of (number density of ).
With a nova system mimicking LMCN 1971-08a in a prepopulated ISM estimated for the local region, it would take years over 62,400 eruptions to reach a radius of 100 pc, containing an evacuated cavity region 15 pc in radius. The NSR outer shell would have a mass of (made up almost exclusively of ISM) and an expansion velocity of , which is consistent with the expansion velocity derived from the Hi position-velocity analysis from Section 3.1.
4 Discussion
4.1 Comparison to other NSRs
With a diameter of pc, the LMCN 1971-08a NSR is larger than the NSR around M31N 2008-12a (major axis of 134 pc; Darnley et al., 2019) and a factor of , and larger than the NSRs associated with KT Eri (50 pc; Shara et al., 2024a), T CrB (30 pc; Shara et al., 2024c) and RS Oph (70 pc; Shara et al., 2025), respectively. The projected morphology of the NSR is also more circular than the other NSRs, likely resulting from the system’s inclination, and is most alike the M31N 2008-12a NSR with more defined edges than the irregular morphology of the Galactic NSRs.
As with all other NSRs, this NSR is most prominent through its H and emission. The H surface brightnesses of the brightest part of the NSR (northeast component) is . Having a H surface brightness comparable to the other NSRs: for KT Eri (Shara et al., 2024a), for T CrB (Shara et al., 2024c), for RS Oph (Shara et al., 2025) and for M31N 2008-12a (based on integrated H flux; Darnley et al. 2019), this NSR is intrinsically brighter than all other NSRs, supporting predictions that larger NSRs should be brighter (Healy-Kalesh et al., 2023). Moreover, the NSR shell is almost negligible through the narrowband filter (Figure 7) due to the lack of photoionisation across the structure, as found for the other NSRs (Darnley et al., 2019; Shara et al., 2024a, c, 2025), though the southwest component does have very faint emission coinciding with a higher /H ratio.
The Hi position-velocity analysis indicates a possible expansion velocity of the NSR outer edges to be . These velocities are consistent with those expansion velocities seen in other NSRs shells (Darnley et al., 2019; Shara et al., 2024a, 2025) and fits with the range of NSR outer shell velocities predicted from modelling in this work and other studies (e.g. Healy-Kalesh et al., 2023).
Both LMCN 1971-08a and another NSR-hosting nova KT Eri show ”several remarkable similarities” (Bode et al., 2016). Those authors suggested both novae may belong to a class of RNe with WDs approximately away from the Chandrasekhar limit but with very high mass transfer rates between eruptions. This similarity may indicate a potential correlation between these types of novae and the brightest NSRs, but cannot be tested further until the sample of NSRs increases.
4.2 Other origins
It is necessary to consider other origins of the visible shell encompassing LMCN 1971-08a including the structure being a supernova remnant (SNR), a superbubble, a planetary nebula (PN) or a Hii region. We therefore searched for sources within the NSR ( radius) using SIMBAD (Wenger et al., 2000) and derived emission line ratios (/H, /H and /[S ii]; see Section A.3) for consideration of other possible scenarios.
While the structure is bright in H and , the low levels of emission (see Section A.3 and Figure 7) and lack of radio emission (see MeerKAT panel in Figure 1) from the shell (both typical of SNR shock fronts) combined with zero radio sources found within the structure, low expansion velocities inferred from Hi data (see Section 3.1) and it being larger than any SNR in the LMC (e.g. Zangrandi et al., 2024) demonstrate that the discovered structure is not a SNR. The absence of strong emission (see Figure 7) and the size of the shell (LMC PNe are unresolved in MCELS data as they are arcsec-scale; e.g. Shaw et al., 2001) rules out a PN origin. There is one young stellar object (Whitney et al., 2008), one cluster of stars (Bica et al., 2008) and two associations of stars (Bica et al., 2008) within the shell but their classification does not elaborate as to whether they are OB associations (Bica et al., 2008), which could be responsible for ionising the local ISM and growing a bubble of this size (Wright, 2020). Also, the structure is not included in the Hii region catalogue by Pellegrini et al. (2012) so rules against it being a Hii region or a superbubble, as does the lack of radio emission and relatively high /H ratio (see Figure 8).
4.3 Confirmation of recurrent nature
LMCN 1971-08a is deemed a recurrent nova on account of its spatial coincidence with the location of nova N LMC 1971b (Bode et al., 2016). Though possibly some skepticism remained about the recurrent nature of the nova; for example, the entry for LMC 2009-02 (another identifier for LMCN 1971-08a) in Table 7 of Orio et al. (2021) is labelled as ‘perhaps’ a RN, unlike other accepted RNe in the LMC. Therefore, the discovery of a NSR around this nova further supports the status of an observationally-defined RN, but from a different approach.
In addition, owing to the existence of such a NSR, LMCN 1971-08a may have a much shorter recurrence period than the 38 years currently presumed and may therefore be observed in eruption much sooner than 2047.
5 Summary
In this Letter, we detailed the discovery of a nova super-remnant in the LMC hosted by the recurrent nova LMCN 1971-08a. The structure is spatially coincident with the location of the nova, is 200 pc in diameter making it the largest NSR found and is visible in narrowband H and emission indicative of shock-ionisation but with negligible emission. A corresponding structure is apparent in Hi data. Joining the growing collection of such NSRs, this NSR is the first to be discovered in the LMC, the brightest NSR to date and only the second extragalactic nova shell found after the M31 NSR (Darnley et al., 2019).
Acknowledgements.
We thank the referee for feedback that improved our study. MS and FZ acknowledges support from the DFG through the grants SA 2131/13-2, SA 2131/14-2, and SA 2131/15-2. This work made use of the high performance computing facilities at Liverpool John Moores University, partly funded by LJMU’s FET and by the Royal Society. This project used data obtained with the Dark Energy Camera (DECam), which was constructed by the Dark Energy Survey (DES) collaboration. Funding for the DES Projects has been provided by the DOE and NSF (USA), MISE (Spain), STFC (UK), HEFCE (UK), NCSA (UIUC), KICP (U. Chicago), CCAPP (Ohio State), MIFPA (Texas A&M), CNPQ, FAPERJ, FINEP (Brazil), MINECO (Spain), DFG (Germany) and the Collaborating Institutions in the Dark Energy Survey, which are Argonne Lab, UC Santa Cruz, University of Cambridge, CIEMAT-Madrid, University of Chicago, University College London, DES-Brazil Consortium, University of Edinburgh, ETH Zürich, Fermilab, University of Illinois, ICE (IEEC-CSIC), IFAE Barcelona, Lawrence Berkeley Lab, LMU München and the associated Excellence Cluster Universe, University of Michigan, NSF NOIRLab, University of Nottingham, Ohio State University, OzDES Membership Consortium, University of Pennsylvania, University of Portsmouth, SLAC National Lab, Stanford University, University of Sussex, and Texas A&M University. Based on observations at NSF Cerro Tololo Inter-American Observatory, NSF NOIRLab (NOIRLab Prop. ID 2018B-0908; PI: T. Puzia) and (NOIRLab Prop. ID 2021B-0060; PI: S. Points), which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the U.S. National Science Foundation. This research has made use of the SIMBAD database, operated at CDS, Strasbourg, France (see Wenger et al., 2000). This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France. The original description of the VizieR service was published in Ochsenbein et al. (2000).References
- Bica et al. (2008) Bica, E., Bonatto, C., Dutra, C. M., & Santos, J. F. C. 2008, MNRAS, 389, 678
- Bode et al. (2016) Bode, M. F., Darnley, M. J., Beardmore, A. P., et al. 2016, ApJ, 818, 145
- Caldwell et al. (2025) Caldwell, N., Raymond, J. C., Long, K. S., & Lee, M. G. 2025, ApJ, 983, 150
- Darnley et al. (2019) Darnley, M. J., Hounsell, R., O’Brien, T. J., et al. 2019, Nature, 565, 460
- Flaugher et al. (2015) Flaugher, B., Diehl, H. T., Honscheid, K., et al. 2015, AJ, 150, 150
- Graham (1971) Graham, J. A. 1971, IAU Circ., 2353, 1
- Guerrero et al. (2013) Guerrero, M. A., Toalá, J. A., Medina, J. J., et al. 2013, A&A, 557, A121
- Healy-Kalesh (2024) Healy-Kalesh, M. 2024, in The Golden Age of Cataclysmic Variables and Related Objects - VI, 35
- Healy-Kalesh et al. (2023) Healy-Kalesh, M. W., Darnley, M. J., Harvey, É. J., et al. 2023, MNRAS, 521, 3004
- Healy-Kalesh et al. (2024a) Healy-Kalesh, M. W., Darnley, M. J., Harvey, É. J., & Newsam, A. M. 2024a, MNRAS, 529, L175
- Healy-Kalesh et al. (2024b) Healy-Kalesh, M. W., Darnley, M. J., & Shara, M. M. 2024b, MNRAS, 528, 3531
- Healy-Kalesh et al. (2024c) Healy-Kalesh, M. W., Darnley, M. J., Shara, M. M., et al. 2024c, MNRAS, 529, 236
- HI4PI Collaboration et al. (2016) HI4PI Collaboration, Ben Bekhti, N., Flöer, L., et al. 2016, A&A, 594, A116
- Hillman et al. (2016) Hillman, Y., Prialnik, D., Kovetz, A., & Shara, M. M. 2016, ApJ, 819, 168
- Jonas & MeerKAT Team (2016) Jonas, J. & MeerKAT Team. 2016, in MeerKAT Science: On the Pathway to the SKA, 1
- Kerp et al. (2011) Kerp, J., Winkel, B., Ben Bekhti, N., Flöer, L., & Kalberla, P. M. W. 2011, Astronomische Nachrichten, 332, 637
- Lanzetta et al. (2023) Lanzetta, K. M., Gromoll, S., Shara, M. M., et al. 2023, PASP, 135, 015002
- Liller (2009) Liller, W. 2009, IAU Circ., 9019, 1
- McClure-Griffiths et al. (2009) McClure-Griffiths, N. M., Pisano, D. J., Calabretta, M. R., et al. 2009, ApJS, 181, 398
- O’Brien et al. (2001) O’Brien, T. J., Davis, R. J., Bode, M. F., Eyres, S. P. S., & Porter, J. M. 2001, in International Astronomical Union, Vol. 205, Galaxies and their Constituents at the Highest Angular Resolutions, ed. R. T. Schilizzi, 260
- Ochsenbein et al. (2000) Ochsenbein, F., Bauer, P., & Marcout, J. 2000, A&AS, 143, 23
- Orio et al. (2021) Orio, M., Dobrotka, A., Pinto, C., et al. 2021, MNRAS, 505, 3113
- Orio et al. (2009) Orio, M., Mason, E., Gallagher, J., & Abbott, T. 2009, The Astronomer’s Telegram, 1930, 1
- Peimbert et al. (2017) Peimbert, M., Peimbert, A., & Delgado-Inglada, G. 2017, PASP, 129, 082001
- Pellegrini et al. (2012) Pellegrini, E. W., Oey, M. S., Winkler, P. F., et al. 2012, ApJ, 755, 40
- Pietrzyński et al. (2019) Pietrzyński, G., Graczyk, D., Gallenne, A., et al. 2019, Nature, 567, 200
- Points et al. (2024) Points, S. D., Long, K. S., Blair, W. P., et al. 2024, ApJ, 974, 70
- Santamaría et al. (2025) Santamaría, E., Guerrero, M. A., Ramos-Larios, G., Toalá, J. A., & Sabin, L. 2025, MNRAS, 539, 246
- Sasaki et al. (2025) Sasaki, M., Zangrandi, F., Filipović, M., et al. 2025, A&A, 693, L15
- Shara et al. (2024a) Shara, M. M., Lanzetta, K. M., Garland, J. T., et al. 2024a, MNRAS, 529, 224
- Shara et al. (2024b) Shara, M. M., Lanzetta, K. M., Garland, J. T., et al. 2024b, MNRAS, 529, 212
- Shara et al. (2025) Shara, M. M., Lanzetta, K. M., Masegian, A., et al. 2025, AJ, 170, 56
- Shara et al. (2024c) Shara, M. M., Lanzetta, K. M., Masegian, A., et al. 2024c, ApJ, 977, L48
- Shaw et al. (2001) Shaw, R. A., Stanghellini, L., Mutchler, M., Balick, B., & Blades, J. C. 2001, ApJ, 548, 727
- Smeaton et al. (2025) Smeaton, Z. J., Filipović, M. D., Alsaberi, R. Z. E., et al. 2025, Serbian Astronomical Journal, 210, 13
- Smith & Team (1999) Smith, R. C. & Team, M. 1999, in IAU Symp., Vol. 190, 28
- Starrfield et al. (1972) Starrfield, S., Truran, J. W., Sparks, W. M., & Kutter, G. S. 1972, ApJ, 176, 169
- Subramanian & Subramaniam (2009) Subramanian, S. & Subramaniam, A. 2009, A&A, 496, 399
- Tosi et al. (2024) Tosi, S., Dell’Agli, F., Kamath, D., et al. 2024, A&A, 688, A36
- Vaytet et al. (2007) Vaytet, N. M. H., O’Brien, T. J., & Bode, M. F. 2007, ApJ, 665, 654
- Wenger et al. (2000) Wenger, M., Ochsenbein, F., Egret, D., et al. 2000, A&AS, 143, 9
- Whelan & Iben (1973) Whelan, J. & Iben, Icko, J. 1973, ApJ, 186, 1007
- Whitney et al. (2008) Whitney, B. A., Sewilo, M., Indebetouw, R., et al. 2008, AJ, 136, 18
- Worters et al. (2007) Worters, H. L., Eyres, S. P. S., Bromage, G. E., & Osborne, J. P. 2007, MNRAS, 379, 1557
- Wright (2020) Wright, N. J. 2020, New A Rev., 90, 101549
- Yaron et al. (2005) Yaron, O., Prialnik, D., Shara, M. M., & Kovetz, A. 2005, ApJ, 623, 398
- Yew et al. (2021) Yew, M., Filipović, M. D., Stupar, M., et al. 2021, MNRAS, 500, 2336
- Zangrandi et al. (2024) Zangrandi, F., Jurk, K., Sasaki, M., et al. 2024, A&A, 692, A237
Appendix A Further remarks
A.1 NSR missed in previous work
The LMC was surveyed in previous work (Healy-Kalesh et al. 2024b) through a targeted search around the four recurrent novae using the Faulkes Telescope South but no NSRs were uncovered. While arcminute FOV images were provided in Healy-Kalesh et al. (2024b) around LMCN 1971-08a, the field of view of the Faulkes Telescope South facility used was arcminutes (chosen to search the surroundings of the RNe for this reason by Healy-Kalesh et al. (2024b)). As such, the region observed was large enough to capture a component of the NSR, however the relatively shorter exposure time of the observations (s for H) and therefore fainter emission, in addition to (and more importantly) only a component of the NSR being captured, indicates why this structure was overlooked in the previous work. The FOV of the Faulkes Telescope South observation from Healy-Kalesh et al. (2024b) is shown in Figure 6. Now with knowledge of the gross structure of the NSR, the southwest shell edge can be made out when comparing with the DeMCELS image in Figure 6, as can the structure to the west of the nova’s position.
A.2 Surroundings of other LMC recurrent novae
The surroundings of all four RNe in the LMC were initially checked within the H data of MCELS for signs of a NSR. While the NSR was discovered around LMCN 1971-08a, little is seen in the environments around LMCN 1968-12a, LMCN 1996 and YY Doradus.
A.3 Comparison of line ratios with other possible structures
The ratios of emission lines can be used as diagnostic tools for helping to differentiate between different extended structures. As such, in Figure 7, we compare the ratio for the nova super-remnant with the same ratio for three LMC supernova remnants, J0450-7050 (a.k.a. Veliki; Smeaton et al. 2025) and J0534-7033 and J0506-6541 (from Zangrandi et al. 2024); the LMC Hii region, L77 DEM L45 (from Pellegrini et al. 2012); and the LMC planetary nebula, SMP LMC 80 (from Tosi et al. 2024), in order to differentiate the shell from other structures. We also provide similar comparative plots for the ratios and in Figure 8 and Figure 9, respectively.
The ratios were calculated for a number of apertures (with radii of ) along a slit passing through the structure (illustrated in Figures 7, 8 and 9). As the structures are different sizes, the length of the line (and therefore the number of apertures used) were scaled accordingly such that the ratios can be compared. The slit shown in Figure 7 is long with 100 apertures for the nova super-remnant; long with 4 apertures for planetary nebula SMP LMC 80 (as PNe in the LMC are compact); long with 80 apertures for Hii region L77 DEM L45; long with 40 apertures for SNR J0450-7050; long with 25 apertures for SNR J0534-7033; and long with 50 apertures for SNR J0506-6541. In the ratio plots of Figures 7, 8 and 9, the axes have been normalised.
Supernova remnants:
The emission line ratio for supernova remnants is typically high ( and see e.g. Yew et al. 2021; Caldwell et al. 2025) as a result of enhanced ionisation and excitation of sulphur from radiative shocks (Zangrandi et al. 2024). In addition to this, SNRs also exhibit strong emission as a result of shock heating (Yew et al. 2021). While the nova super-remnant does exhibit a ratio on par with the example supernova remnants ( and see Figure 8), there is a large difference in the ratios for the NSR and the SNRs, with the NSR emission being much fainter () within the NE and SW component of the shell (between and on the axis of Figure 7).
Planetary nebulae:
Planetary nebulae have bright emission (Guerrero et al. 2013, and see the example PN in Figure 7) as a result of fast outflows shocking the surrounding low-density material (Guerrero et al. 2013) and weak emission (see, e.g. Caldwell et al. 2025). However, as can be seen in Figure 7, the nova super-remnant is faint in (with in the NE and SW component of the shell) and has moderate levels of emission (see Figure 8).
Hii regions:
Typical Hii regions have low ratios as a result of the sulphur in such structures existing primarily as doubly-ionised sulphur (Yew et al. 2021) from the gas being photoionised (Zangrandi et al. 2024). As can be seen in Figure 8, the nova super-remnant has a higher ratio across the shell than the ratio of a typical Hii region. In addition, as illustrated with the example Hii region in Figure 7, Hii regions have high levels of due to photoionisation from hot central stars (Peimbert et al. 2017) however the discovered structure does not with for the NSR compared to for the Hii region.
Appendix B Figures